21 September 2011

Until 09:40, work done on MODS restart.
  • Guider camera swapped out in the afternoon.
  • Star images on new Guider camera smeared to +X: aberrations or camera. Seems to be camera or DSP SW. Will be investigated in afternoon.
  • Despite smeared images, initial settings for ROI, guide camera center and hotspot determined.
  • To work on secondary, tried to switch to LFBG and IRTC, but left front rotator problem.
  • Later, jet problems.
  • Finally left rotator running, but AGW problems.
  • 09:40, switch to LBCs.

09:15 Some discussion with John Hill about image quality on Blue. In the evening I put some images from superfoc sequences and extra/intra focal sequences at http://ohia.as.arizona.edu/~kuhn/LBCimages20110920/. There is a skirt on the blue pupils which appears to be a rotational trail. He notes CMU computer is 10 seconds off. This needs to be corrected. Incorrect time could cause incorrect track rate and/or rotator to stop tracking before exposure is completed. The resulting trails would be misinterpreted as astigmatism.

09:45 Plan to observe OSURC science targets (N672 and 628). Also to obtain SDT and B flats in morning twilight.

10:00 Slew to WT10_280. Focus/collimate with dofpia,/first. Observe and run lbcrangebal.

10:20 Slew to focus field for N672. Run dofpia.

10:27 Run N672 monitor OB.

about 11:00 Run FPIA again on focus field and obtain 2 300-sec exposure at U-BESSEL (because of problem with SDT_Uspec filter).

11:22 Slew to N628 focus field.

11:31 Start N628 OB. Copy and edit OB to do U-BESSEL also. Initial images (2xB and 2xR) are elongated. Guide star on Red is elongated. Pause OB.

11:43 refocus/collimate. Edit U-BESSEL version to remove B-BESSEL and do V, S and U. Red guide star is again elongated - it is not just one, but all guide stars on the red chip look this way. Tech chip is out of focal plane.. images before looked better than guide star, but still elongated. Elevation 44 degrees.

12:09 Issue trak 3664 submitted regarding elongated tech chip images (on some fields).

12:17 Stopped OB after first pair of U-Bessel/R-Bessel - getting light.

12:17 Norm changed time on CMU. It was 14 sec slow. I slew to field higher than the one we were observing (which was at elev 44). Field about 75 deg.

12:22 dofpia on this field at 74 deg elevation. Converged in 2 for B but red taking longer (first z4 was -2000 nm). For Blue, corrections from final pupil images:
   141    -25     42    -61    -31    -43     10    y  0.89  0.11                      /newdata/lbcb.20110921.122603.fits  V-BESSEL

12:27 exin4 to take extra/intra focal pair to see whether "skirt" around Blue pupil images is field dependent and suggests rotational trailing during the 16-sec exposures.
  • 122745/122741 are extra (hmmm - hard to see "skirt". I do see "coma" or illumination of one part of pupil more than other and this is field dependent - brighter side seems to be roughly towards center).

12:37 Flats:
  • Doing Bz flats 4-sec 30K
  • SDT_Uspec/Y: Obtained set of PA=0 Uspec/Y flats and then one pair at PA=180: 124817 is SDT_Uspec flat at PA=180.
  • Note that flats through SDT_Uspec were obtained. These did not look abnormal. The cause of the haloes around bright stars through this filter is under investigation (IT #3655).

12:50 A set of biases (but while closing).

Record of Z4 corrections sent (note that these two OBs were created before tech chip AF and may not have had stars on the Active Optics tech chip (tech #2).
2011/09/21 10:56:25.006455 S R TRACKERS          Active Optics P03 -2.00000000 +0.00000000
2011/09/21 10:56:41.036236 S B TRACKERS          Active Optics P03 +0.00000000 +0.00000000
2011/09/21 11:02:03.773011 S R TRACKERS          Active Optics P03 -153.00000000 +0.00000000
2011/09/21 11:02:24.919295 S B TRACKERS          Active Optics P03 +0.00000000 +0.00000000
2011/09/21 11:16:29.391273 S R TRACKERS          Active Optics P03 +0.00000000 +0.00000000
2011/09/21 11:16:33.247538 S B TRACKERS          Active Optics P03 +0.00000000 +0.00000000
2011/09/21 11:22:15.597450 S B TRACKERS          Active Optics P03 +0.00000000 +0.00000000
2011/09/21 11:37:04.247234 S R TRACKERS          Active Optics P03 +0.00000000 +0.00000000
2011/09/21 11:37:06.996540 S B TRACKERS          Active Optics P03 +0.00000000 +0.00000000
2011/09/21 11:42:43.290418 S R TRACKERS          Active Optics P03 +203.00000000 +0.00000000
2011/09/21 11:42:50.157028 S B TRACKERS          Active Optics P03 +0.00000000 +0.00000000
2011/09/21 11:55:34.713944 S R TRACKERS          Active Optics P03 +0.00000000 +0.00000000
2011/09/21 11:55:48.487022 S B TRACKERS          Active Optics P03 +0.00000000 +0.00000000

14:00 Another set of biases, but with lights on in dome.

-- OlgaKuhn - 21 Sep 2011
Topic revision: r4 - 22 Sep 2011, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback